OpenStudy (ttop0816):

Use the rules of logarithms to expand each of the following. q: log4 (4x/y^3) (my answer): log4 (4x/y^3) log4 (4x) - log4 (y^3) log4 (4x) - 3log4 (y) Am I correct??

OpenStudy (dayakar):

C is correct

OpenStudy (dayakar):

log m/n = logm-log n

OpenStudy (dayakar):

log a^m = m log a

OpenStudy (dayakar):

do u get it u want explanation

OpenStudy (ttop0816):

Nono I just wrote down how I got my answer, its not multiple choice haha